OpenStudy (anonymous):

Solve the triangle. A = 54°, b = 11, c = 8 a ≈ 13.4, C ≈ 45.6, B ≈ 80.4 No triangles possible a ≈ 9, C ≈ 45.6, B ≈ 80.4 a ≈ 13.4, C ≈ 49.6, B ≈ 76.4

OpenStudy (anonymous):

need the law of cosines for this one

OpenStudy (anonymous):

\[a^2=b^2+c^2-2ab\cos(A)\] \[a^2=11^2+8^2-2\times 11\times 8\cos(54)\] and a calculator

OpenStudy (anonymous):

typo in the first line, should have been \[a^2=b^2+c^2-2bc\cos(A)\] second line is right though, and that is what you need